# Learning Unsigned Distance Fields from Local Shape Functions for 3D Surface Reconstruction

### v1

@@ Abstract
Unsigned distance fields (UDFs) provide a versatile framework for representing a diverse array of 3D shapes, encompassing both watertight and non-watertight geometries. Traditional UDF learning methods typically require extensive training on large datasets of 3D shapes, which is costly and often necessitates hyperparameter adjustments for new datasets. This paper presents a novel neural framework, LoSF-UDF, for reconstructing surfaces from 3D point clouds by leveraging local shape functions to learn UDFs. We observe that 3D shapes manifest simple patterns within localized areas, prompting us to create a training dataset of point cloud patches characterized by mathematical functions that represent a continuum from smooth surfaces to sharp edges and corners. Our approach learns features within a specific radius around each query point and utilizes an attention mechanism to focus on the crucial features for UDF estimation. This method enables efficient and robust surface reconstruction from point clouds without the need for shape-specific training. Additionally, our method exhibits enhanced resilience to noise and outliers in point clouds compared to existing methods. We present comprehensive experiments and comparisons across various datasets, including synthetic and real-scanned point clouds, to validate our method’s efficacy.
@@ 1 Introduction
In this paper, we propose a lightweight and effective supervised learning framework, Losf-UDF, to address these challenges. Since learning UDFs does not require determining whether a query point is inside or outside the geometry, it is a local quantity independent of the global context. Inspired by the observation that 3D shapes manifest simple patterns within localized areas, we synthesize a training dataset comprising a set of point cloud patches by utilizing local shape functions. Subsequently, we can estimate the unsigned distance values by learning local geometric features through an attention-based network. Our approach distinguishes itself from existing methods by its novel training strategy. Specifically, it is uniquely trained on synthetic surfaces, yet it demonstrates remarkable capability in predicting UDFs for a wide range of common surface types. For smooth surfaces, we generate training patches (quadratic surfaces) by analyzing principal curvatures, meanwhile, we design simple shape functions to simulate sharp features. This strategy has three unique advantages. First, it systematically captures the local geometries of most common surfaces encountered during testing, effectively mitigating the dataset dependence risk that plagues current UDF learning methods. Second, for each training patch, the ground-truth UDF is readily available, streamlining the training process. Third, this approach substantially reduces the costs associated with preparing the training datasets. We evaluate our framework on various datasets and demonstrates its ability to robustly reconstruct high-quality surfaces, even for point clouds with noise and outliers. Notably, our method can serve as a lightweight initialization that can be integrated with existing unsupervised methods to enhance their performance. We summarize our main contributions as follows.
@@
  • Our framework achieves superior performance in surface reconstruction from both synthetic point clouds and real scans, even in the presence of noise and outliers.
@@ 2 Related Work
Surface reconstruction. Reconstructing 3D surfaces from point clouds is a classic and important topic in computer graphics. The most widely used Poisson method [1, 22] fits surfaces by solving PDEs. These traditional methods involve adjusting the gradient of an indicator function to align with a solution derived from a (screened) Poisson equation. A crucial requirement of these methods is the input of oriented normals. The Iterative Screened Poisson Reconstruction method[23] introduced an iterative approach to refine the reconstruction process, improving the ability to generate surfaces from point clouds without direct computation of normals. The shape of points [24] introduced a differentiable point-to-mesh layer by employing a differentiable formulation of PSR [1] to generate watertight, topology-agnostic manifold surfaces.
@@
Neural surface representations. Recently, the domain of deep learning has spurred significant advances in the implicit neural representation of 3D shapes. Some of these works trained a classifier neural network to construct occupancy fields [9, 10, 11, 12] for representing 3D geometries. Poco [12] achieves superior reconstruction performance by introducing convolution into occupancy fields. Ouasfi et al. [25] recently proposed a uncertainty measure method based on margin to learn occupancy fields from sparse point clouds. Compared to occupancy fields, SDFs [2, 3, 4, 5, 6, 7, 8] offer a more precise geometric representation by differentiating between interior and exterior spaces through the assignment of signs to distances. Some recent SOTA methods, such as DeepLS [3], using volumetric SDFs to locally learned continuous SDFs, have achieved higher compression, accuracy, and local shape refinement.
@@
Unsigned distance fields learning. Although Occupancy fields and SDFs have undergone significant development recently, they are hard to reconstruct surfaces with boundaries or nonmanifold features. G-Shell[26] developed a differentiable shell-based representation for both watertight and non-watertight surfaces. However, UDFs provide a simpler and more natural way to represent general shapes [13, 14, 15, 16, 17, 18, 19, 20]. Various methods have been proposed to reconstruct surfaces from point clouds by learning UDFs. CAP-UDF [17] suggested directing 3D query points towards the surface with a consistency constraint to develop UDFs that are aware of consistency. LevelSetUDF [14] learned a smooth zero-level function within UDFs through level set projections. As a supervised approach, GeoUDF [15] estimates UDFs by learning local geometric priors from training on many 3D shapes. DUDF [19] formulated the UDF learning as an Eikonal problem with distinct boundary conditions. UODF [20] proposed unsigned orthogonal distance fields that every point in this field can access to the closest surface points along three orthogonal directions. Instead of reconstructing from point clouds, many recent works [27, 28, 29, 30] learn high-quality UDFs from multi-view images for reconstructing non-watertight surfaces. Furthermore, UiDFF [31] presents a 3D diffusion model for UDFs to generate textured 3D shapes with boundaries.
@@ 3 Method
Motivation. Distinct from SDFs, there is no need for UDFs to determine the sign to distinguish between the inside and outside of a shape. Consequently, the UDF values are solely related to the local geometric characteristics of 3D shapes. Furthermore, within a certain radius for a query point, local geometry can be approximated by general mathematical functions. Stemming from these insights, we propose a novel UDF learning framework that focuses on local geometries. We employ local shape functions to construct a series of point cloud patches as our training dataset, which includes common smooth and sharp geometric features. Fig. 1 illustrates the pipeline of our proposed UDF learning framework.
@@ 3.1 Local shape functions
$[\mathrm{I}]_{\mathbf{p}}=\begin{bmatrix}E&F\\ F&G\end{bmatrix}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad[\mathrm{II}]_{\mathbf{p}}=\begin{bmatrix}L&M\\ M&N\end{bmatrix}=\begin{bmatrix}\kappa_{1}&0\\ 0&\kappa_{2}\end{bmatrix},\quad$ (1)
@@
$\begin{split}\mathbf{r}(u,v)&=\mathbf{r}(0,0)+\mathbf{r}_{u}(0,0)u+\mathbf{r}_{v}(0,0)v+\frac{1}{2}[\mathbf{r}_{uu}(0,0)u^{2}+\\ &\mathbf{r}_{uv}(0,0)uv+\mathbf{r}_{vv}(0,0)v^{2}]+o(u^{2}+v^{2}).\end{split}$ (2)
Decomposing $\mathbf{r}_{uu}(0,0),\mathbf{r}_{uv}(0,0),\text{and }\mathbf{r}_{vv}(0,0)$ along the tangential and normal directions, we can formulate Eq.(2) according to Eq.(1) as
$\begin{split}\mathbf{r}(u,v)&=\mathbf{r}(0,0)+(u+o(\sqrt{u^{2}+v^{2}}))\mathbf{e}_{1}+(v+o(\sqrt{u^{2}+v^{2}}))\mathbf{e}_{2}\\ &+\frac{1}{2}(\kappa_{1}u^{2}+\kappa_{2}v^{2}+o(u^{2}+v^{2})))\mathbf{n}\end{split}$ (3)
@@
(a) Smooth patches (b) Sharp patches
Figure 3: Synthetic dataset for training. By manipulating functional parameters, we can readily create various smooth and sharp surfaces, subsequently acquiring pairs of point cloud patches and query points via sampling.
Sharp patches. For surfaces with sharp features, they are not differentiable at some points and cannot be approximated in the form of a quadratic surface. We categorize commonly seen sharp geometric features into four types, including creases, cusps, corners, and v-saddles, as illustrated in Fig. 2(b). We construct these four types of sharp features in a consistent form $z=f(x,y)$ like smooth patches
$\begin{split}&\text{creases:}\,z=1-h\cdot\frac{|kx-y|}{\sqrt{1+k^{2}}},~{}\text{cusps:}\,z=1-h\cdot\sqrt{x^{2}+y^{2}},\\ &\text{corners:}\,z=1-h\cdot\max(|x|,|y|),~{}\text{v-saddles:}\,z=1-h\cdot|x|+|y|\cdot(\frac{|x|}{x}\cdot\frac{|y|}{y}),\\ \end{split}$ (5)
where $h$ can adjust the sharpness of the shape, and $k$ can control the direction of the crease. Fig 3 illustrates various smooth and sharp patches with distinct parameters.
Synthetic training dataset. We utilize the mathematical functions introduced above to synthesize a series of point cloud patches for training. As shown in Fig. 3, we first uniformly sample $m$ points $\{(x_{i},y_{i})\}_{i=1}^{m}$ within a circle of radius $r_{0}$ centered at $(0,0)$ in the $xy$-plane. Then, we substitute the coordinates into Eq.(4-5) to obtain the corresponding $z$-coordinate values, resulting in a patch $\mathcal{P}=\{\mathbf{p}_{i=1}^{m}\}$, where $\mathbf{p}_{i}=(x_{i},y_{i},z(x_{i},y_{i}))$. Subsequently, we randomly collect query points $\{\mathbf{q_{i}}\}_{i=1}^{n}$ distributed along the vertical ray intersecting the $xy$-plane at the origin, extending up to a distance of $r_{0}$. For each query point $\mathbf{q}_{i}$, we determine its UDF value $\mathcal{U}(\mathbf{q}_{i})$, which is either $|\mathbf{q}_{i}^{(z)}|$ for smooth patches or $1-|\mathbf{q}_{i}^{(z)}|$ for sharp patches. Noting that for patches with excessively high curvature or sharpness, the minimum distance of the query points may not be the distance to $(0,0,z(0,0))$, we will exclude these patches from our training dataset. Overall, each sample in our synthetic dataset is specifically in the form of $\{\mathbf{q},\mathcal{P},\mathcal{U}(\mathbf{q})\}$.
@@ 3.2.1 Network architecture
where we take $\mathbf{f}_{p}$ as the Key-Value (KV) pair and $\mathbf{f}_{v}$ as the Query (Q). In our experiments, we set $l_{p}=l_{v}=64$, and $l_{G}=128$. Based on the learned geometric features, we aim to fit the UDF values from the distance within the local point cloud. Therefore, we concatenate the distances $\mathbf{d}\in\mathbb{R}^{m}$ induced from $\mathcal{V}$ with the latent code $\mathbf{f}_{G}$, followed by a series of fully connected layers to output the predicted UDF values $\mathcal{U}_{\Theta}(\mathbf{q})$. Fig. 4 illustrates the overall network architecture and data flow. The two PointNets used in our network to extract features from point cloud patches $\mathcal{P}$ and vectors $\mathcal{V}$ consist of four ResNet blocks. In addition, the two fully connected layer modules in our framework consist of three layers each. To ensure non-negativity of the UDF values output by the network, we employ the softplus activation function.
@@ 3.2.2 Training and evaluation
Loss functions. We employ $L_{1}$ loss $\mathcal{L}_{\text{u}}$ to measure the discrepancy between the predicted UDF values and the GT UDF values. Moreover, for the displacements $\Delta\mathbf{d}$ output by the denoising module, we employ $L_{1}$ regularization to encourage sparsity. Consequently, we train the network driven by the following loss function,
$\mathcal{L}=\mathcal{L}_{u}+\lambda_{d}\mathcal{L}_{r},\quad\text{where}\,\,\mathcal{L}_{u}=|\mathcal{U}(\mathbf{q})-\mathcal{U}_{\Theta}(\mathbf{q})|,\,\,\mathcal{L}_{r}=|\Delta\mathbf{d}|,$ (7)
where we set $\lambda_{d}=0.01$ in our experiments.
@@
4 Experiments
4.1 Experiment setup

### v2

# A Lightweight UDF Learning Framework for 3D Reconstruction Based on Local Shape Functions

@@ Abstract
Unsigned distance fields (UDFs) provide a versatile framework for representing a diverse array of 3D shapes, encompassing both watertight and non-watertight geometries. Traditional UDF learning methods typically require extensive training on large 3D shape datasets, which is costly and necessitates re-training for new datasets. This paper presents a novel neural framework, LoSF-UDF, for reconstructing surfaces from 3D point clouds by leveraging local shape functions to learn UDFs. We observe that 3D shapes manifest simple patterns in localized regions, prompting us to develop a training dataset of point cloud patches characterized by mathematical functions that represent a continuum from smooth surfaces to sharp edges and corners. Our approach learns features within a specific radius around each query point and utilizes an attention mechanism to focus on the crucial features for UDF estimation. Despite being highly lightweight, with only 653 KB of trainable parameters and a modest-sized training dataset with 0.5 GB storage, our method enables efficient and robust surface reconstruction from point clouds without requiring for shape-specific training. Furthermore, our method exhibits enhanced resilience to noise and outliers in point clouds compared to existing methods. We conduct comprehensive experiments and comparisons across various datasets, including synthetic and real-scanned point clouds, to validate our method’s efficacy. Notably, our lightweight framework offers rapid and reliable initialization for other unsupervised iterative approaches, improving both the efficiency and accuracy of their reconstructions. Our project and code are available at https://jbhu67.github.io/LoSF-UDF.github.io/.
@@ 1 Introduction
In this paper, we propose a lightweight and effective supervised learning framework, LoSF-UDF, to address these challenges. Since learning UDFs does not require determining whether a query point is inside or outside the geometry, it is a local quantity independent of the global context. Inspired by the observation that 3D shapes manifest simple patterns within localized areas, we synthesize a training dataset comprising a set of point cloud patches by utilizing local shape functions. Subsequently, we can estimate the unsigned distance values by learning local geometric features through an attention-based network. Our approach distinguishes itself from existing methods by its novel training strategy. Specifically, it is uniquely trained on synthetic surfaces, yet it demonstrates remarkable capability in predicting UDFs for a wide range of common surface types. For smooth surfaces, we generate training patches (quadratic surfaces) by analyzing principal curvatures, meanwhile, we design simple shape functions to imitate sharp features. This strategy has three unique advantages. First, it systematically captures the local geometries of most common surfaces encountered during testing, effectively mitigating the dataset dependence risk that plagues current UDF learning methods. Second, for each training patch, the ground-truth UDF is readily available, streamlining the training process. Third, this approach substantially reduces the costs associated with preparing the training datasets. We evaluate our framework on various datasets and demonstrates its ability to robustly reconstruct high-quality surfaces, even for point clouds with noise and outliers. Notably, our method can serve as a lightweight initialization that can be integrated with existing unsupervised methods to enhance their performance. We summarize our main contributions as follows.
@@
  • Our lightweight framework offers rapid and reliable initialization for other unsupervised iterative approaches, improving both efficiency and accuracy.
@@ 2 Related work
Neural surface representations. Reconstructing 3D surfaces from point clouds is a classic and important topic in computer graphics [4, 5, 44]. Recently, the domain of deep learning has spurred significant advances in the implicit neural representation of 3D shapes. Some of these works trained a classifier neural network to construct occupancy fields [29, 10, 35, 6] for representing 3D geometries. Poco [6] achieves superior reconstruction performance by introducing convolution into occupancy fields. Ouasfi et al. [32] recently proposed a uncertainty measure method based on margin to learn occupancy fields from sparse point clouds. Compared to occupancy fields, SDFs [33, 1, 42, 38, 43, 2] offer a more precise geometric representation by differentiating between interior and exterior spaces through the assignment of signs to distances.
@@
Unsigned distance fields learning. Although Occupancy fields and SDFs have undergone significant development recently, they are hard to reconstruct surfaces with boundaries or nonmanifold features. G-Shell [25] developed a differentiable shell-based representation for both watertight and non-watertight surfaces. However, UDFs provide a simpler and more natural way to represent general shapes [11, 49, 37, 46, 48, 22, 15, 27]. Various methods have been proposed to reconstruct surfaces from point clouds by learning UDFs. CAP-UDF [48] suggested directing 3D query points toward the surface with a consistency constraint to develop UDFs that are aware of consistency. LevelSetUDF [49] learned a smooth zero-level function within UDFs through level set projections. As a supervised approach, GeoUDF [37] estimates UDFs by learning local geometric priors from training on many 3D shapes. DUDF [15] formulated the UDF learning as an Eikonal problem with distinct boundary conditions. UODF [27] proposed unsigned orthogonal distance fields that every point in this field can access the closest surface points along three orthogonal directions. Instead of reconstructing from point clouds, many recent works [13, 28, 26, 23] learn high-quality UDFs from multi-view images to reconstruct non-watertight surfaces. Furthermore, UiDFF [50] presents a 3D diffusion model for UDFs to generate textured 3D shapes with boundaries.
Local-based reconstruction. Most methods achieve 3D reconstruction by constructing a global function from point clouds. For example, Poisson methods [19, 20] fits surfaces by solving partial differential equations, while neural network-based methods like DeepSDF [9, 30, 34] represent geometry through network optimization. The limitation of most global methods lies in their need for extensive datasets for training, coupled with inadequate generalization to unseen shape categories. Conversely, 3D surfaces exhibit local similarities and repetitions, which have spurred the development of techniques for reconstructing surfaces locally. Ohtake et al. [31] introduced a shape representation utilizing a multi-scale partition of unity framework, wherein the local shapes of surfaces are characterized by piecewise quadratic functions. DeepLS [7] and LDIF [16] reconstructed local SDF by training learnable implicit functions or neural networks. PatchNets [40] proposed a mid-level patch-based surface representation, facilitating the development of models with enhanced generalizability. Ying et al. [47] introduced a local-to-local shape completion framework that utilized adaptive local basis functions. While these methods all focus on SDF, GeoUDF [37] represents a recent advancement in reconstructing UDF from a local perspective.
@@ 3 Method
Motivation. Distinct from SDFs, there is no need for UDFs to determine the sign to distinguish between the inside and outside of a shape. Consequently, the UDF values are solely related to the local geometric characteristics of 3D shapes. Furthermore, within a certain radius for a query point, local geometry can be approximated by general mathematical functions. Stemming from these insights, we propose a novel UDF learning framework that focuses on local geometries. We employ local shape functions to construct a series of point cloud patches as our training dataset, which includes common smooth and sharp geometric features. Given a point cloud to reconstruct, we employ the optimized model to output the corresponding distance values based on the local patch within radius for each query point. Figure 1 illustrates the pipeline of our proposed UDF learning framework.
@@ 3.1 Local shape functions
$[\mathrm{I}]_{\mathbf{p}}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad[\mathrm{II}]_{\mathbf{p}}=\begin{bmatrix}\kappa_{1}&0\\ 0&\kappa_{2}\end{bmatrix},$ (1)
@@
$\mathbf{r}(u,v)=\mathbf{r}(0,0)+\mathbf{r}_{u}(0,0)u+\mathbf{r}_{v}(0,0)v+\frac{1}{2}[\mathbf{r}_{uu}(0,0)u^{2}$ (2)
$+\mathbf{r}_{uv}(0,0)uv+\mathbf{r}_{vv}(0,0)v^{2}]+o(u^{2}+v^{2}).$
Decomposing $\mathbf{r}_{uu}(0,0),\mathbf{r}_{uv}(0,0),\text{and }\mathbf{r}_{vv}(0,0)$ along the tangential and normal directions, we can formulate Eq. 2 according to Eq. 1 as
$\mathbf{r}(u,v)=\mathbf{r}(0,0)+(u+o(\sqrt{u^{2}+v^{2}}))\mathbf{e}_{1}+(v$ (3)
$+o(\sqrt{u^{2}+v^{2}}))\mathbf{e}_{2}+\frac{1}{2}(\kappa_{1}u^{2}+\kappa_{2}v^{2}+o(u^{2}+v^{2})))\mathbf{n}$
@@
Figure 3: Synthetic surfaces for training. By manipulating functional parameters, we can readily create various smooth and sharp surfaces, subsequently acquiring pairs of point cloud patches and query points via sampling.
Sharp patches. For surfaces with sharp features, they are not differentiable at some points and cannot be approximated in the form of a quadratic surface. We categorize commonly seen sharp geometric features into four types, including creases, cusps, corners, and v-saddles, as illustrated in Fig. 2(b). We construct these four types of sharp features in a consistent form $z=f(x,y)$ like smooth patches. We define a family of functions as
$z=1-h\cdot g(x,y),$ (5)
where $h$ can adjust the sharpness of the shape. Specifically, $g=\frac{|kx-y|}{\sqrt{1+k^{2}}}$ for creases ($k$ can control the direction), $g=\sqrt{x^{2}+y^{2}}$ for cusps, $g=\max(|x|,|y|)$ for corners, $g=(|x|+|y|)\cdot(\frac{|x|}{x}\cdot\frac{|y|}{y})$ for v-saddles. Fig. 3 illustrates several examples of smooth and sharp patches with distinct parameters.
Synthetic training dataset. We utilize the mathematical functions introduced above to synthesize a series of point cloud patches for training. As shown in Fig. 3, we first uniformly sample $m$ points $\{(x_{i},y_{i})\}_{i=1}^{m}$ within a circle of radius $r_{0}$ centered at $(0,0)$ in the $xy$-plane. Then, we substitute the coordinates into Eqs. 4 and 5 to obtain the corresponding $z$-coordinate values, resulting in a patch $\mathcal{P}=\{\mathbf{p}_{i=1}^{m}\}$, where $\mathbf{p}_{i}=(x_{i},y_{i},z(x_{i},y_{i}))$. Subsequently, we randomly collect query points $\{\mathbf{q_{i}}\}_{i=1}^{n}$ distributed along the vertical ray intersecting the $xy$-plane at the origin, extending up to a distance of $r_{0}$. For each query point $\mathbf{q}_{i}$, we determine its UDF value $\mathcal{U}(\mathbf{q}_{i})$, which is either $|\mathbf{q}_{i}^{(z)}|$ for smooth patches or $1-|\mathbf{q}_{i}^{(z)}|$ for sharp patches. Noting that for patches with excessively high curvature or sharpness, the minimum distance of the query points may not be the distance to $(0,0,z(0,0))$, we will exclude these patches from our training dataset. Overall, each sample in our synthetic dataset is specifically in the form of $\{\mathbf{q},\mathcal{P},\mathcal{U}(\mathbf{q})\}$.
@@ 3.2.1 Network architecture
where we take $\mathbf{f}_{p}$ as the Key-Value (KV) pair and $\mathbf{f}_{v}$ as the Query (Q). In our experiments, we set $l_{p}=l_{v}=64$, and $l_{G}=128$. Based on the learned geometric features, we aim to fit the UDF values from the distance within the local point cloud. Therefore, we concatenate the distances $\mathbf{d}\in\mathbb{R}^{m}$ induced from $\mathcal{V}$ with the latent code $\mathbf{f}_{G}$, followed by a series of fully connected layers to output the predicted UDF values $\mathcal{U}_{\Theta}(\mathbf{q})$. Figure 4 illustrates the overall network architecture and data flow. The two PointNets used in our network to extract features from point cloud patches $\mathcal{P}$ and vectors $\mathcal{V}$ consist of four ResNet blocks. In addition, the two fully connected layer modules in our framework consist of three layers each. To ensure non-negativity of the UDF values output by the network, we employ the softplus activation function.
@@ 3.2.2 Training and evaluation
Loss functions. We employ $L_{1}$ loss $\mathcal{L}_{\text{u}}$ to measure the discrepancy between the predicted UDF values and the GT UDF values. Moreover, for the displacements $\Delta\mathbf{d}$ output by the denoising module, we employ $L_{1}$ regularization to encourage sparsity. Consequently, we train the network driven by the loss function $\mathcal{L}=\mathcal{L}_{u}+\lambda_{d}\mathcal{L}_{r}$, where $\mathcal{L}_{u}=|\mathcal{U}(\mathbf{q})-\mathcal{U}_{\Theta}(\mathbf{q})|,\,\,\mathcal{L}_{r}=|\Delta\mathbf{d}|$, we set $\lambda_{d}=0.01$ in our experiments.
@@
3.3 Integration with unsupervised methods
Unsupervised methods, such as CAP-UDF [48] and LevelSetUDF [49], require time-consuming iterative reconstruction of a single point cloud. In contrast, our LoSF-UDF method is a highly lightweight framework. Once trained on a synthetic, shape-independent local patch dataset, it efficiently reconstructs plausible 3D shapes from diverse point clouds, even in the presence of noise and outliers. Although unsupervised methods are time-consuming, they can reconstruct shapes with richer details due to the combined effects of various loss functions. Therefore, we integrate our method with unsupervised approaches to provide better initialization, thereby accelerating convergence (Tab. 3) and achieving improved reconstruction results. Generally, assuming the network of the unsupervised method is $\mathcal{B}_{\Xi}$, we define the loss function of our integrated framework as
$\min\limits_{\Xi}\,\mathcal{L}=\alpha_{t}\frac{1}{N}\sum_{i=1}^{N}|\mathcal{B}_{\Xi}(\mathbf{q}_{i})-\mathcal{U}_{\Theta^{*}}(\mathbf{q}_{i})|+(1-\alpha_{t})\mathcal{L}_{\text{unsupv}},$ (7)
where $\mathcal{B}_{\Xi}$ can be selected as a MLP network like CAP-UDF [48] and LevelSetUDF [49], or a SIREN network [39] like DEUDF [45]. $\mathcal{L}_{\text{unsupv}}$ is the loss functions employed in these unsupervised methods. $\mathcal{U}_{\Theta^{*}}$ is our trained LoSF network with optimized parameters $\Theta^{*}$. $\alpha_{t}\in[0,1]$ is a time-dependent weight. In our experiments (refer to Sec. 4.5), the whole training process requires around 20000 iterations. The value of $\alpha_{t}$ decreases from 1 to 0 gradually during the first 10000 iterations.
4 Experimental results
4.1 Setup